OpenStudy (anonymous):

Which of the following has a slope of –3 and a y-intercept of –4? A. y = –3x + 4 B. y = 3x + 4 C. y = –3x – 4 D. y = 3x – 4 And how do you get the answer for it? I have around 8 more questions like this and I dont understand them

undeadknight26 (undeadknight26):

i spy a trick :D it would be A: - -4 would turn into + 4

OpenStudy (anonymous):

how do you get it though? is there a specific thing for it or something?

OpenStudy (nincompoop):

how?

OpenStudy (nincompoop):

your y-intercept b does not change sign it is as it is

undeadknight26 (undeadknight26):

oh...so it would be...C?

undeadknight26 (undeadknight26):

since y =mx + b m = ur slope and b equals ur y?

OpenStudy (nincompoop):

set x = 0 y = -3(0) - 4 y = -4 in the form of y-intercept form y = mx + b m is your slope and b is your y-intercept

OpenStudy (anonymous):

okay~ thanks